# Sequence Reconstruction for Single-Deletion Single-Substitution Channel

### v1

@@ Abstract
The central problem in sequence reconstruction is to find the minimum number of distinct channel outputs required to uniquely reconstruct the transmitted sequence. According to Levenshtein’s work in 2001, this number is determined by the maximum size of the intersection between the error balls of any two distinct input sequences of the channel. In this work, we study the sequence reconstruction problem for single-deletion single-substitution channel, assuming that the transmitted sequence belongs to a $q$-ary code with minimum Hamming distance at least $2$, where $q\geq 2$ is any fixed integer. Specifically, we prove that for any two $q$-ary sequences with Hamming distance $d\geq 2$, the size of the intersection of their error balls is upper bounded by $2qn-3q-2-\delta_{q,2}$, where $\delta_{i,j}$ is the Kronecker delta. We also prove the tightness of this bound by constructing two sequences the intersection size of whose error balls achieves this bound.
@@ I Introduction
We consider a communication scenario where a codeword $\bm{x}$ from some codebook $\mathscr{C}$ is transmitted over a number of identical channels and the goal is to reconstruct $\bm{x}$ from all (erroneous) channel outputs (also referred to as reads in data storage applications). This problem, also known as the sequence reconstruction problem, was first proposed by Levenshtein [1], [2], and in recent years, gained more and more attentions due to its applications in DNA data storage [3]. The central problem in sequence reconstruction is to determine the minimum number of distinct channel outputs (reads) required to uniquely reconstruct $\bm{x}$. This number was shown to be equal to one plus the maximum size of the intersection between the error balls of any two distinct codewords of $\mathscr{C}$ (also referred to as the read coverage of $\mathscr{C}$ for the channel) [1]. Therefore, deriving the read coverage of $\mathscr{C}$ is essential for solving the sequence reconstruction problem. On the other hand, designing of codes with given read coverage, called reconstruction codes, is also an interesting problem for sequence reconstruction.
@@
In his seminal work [1], Levenshtein studied the sequence reconstruction problem for deletion, insertion, substitution and transposition separately, where $\mathscr{C}$ is taken to be the set of all $q$-ary sequence. For the more general case that $\mathscr{C}$ is an $(\ell-1)$-deletion correcting code for some positive integer $\ell\leq t$, the problem was studied in [4] and [5] for $t$-deletion channel, and in [6] for $t$-insertion channel. Reconstruction codes for two-deletion channels can be found in [7, 8] and reconstruction codes for two-insertion channels can be found in [9]. Reconstruction codes for $q$-ary single-edit channel $(q\geq 2)$ was constructed in [10] by generalizing the construction in [11], where an edit error means a deletion, an insertion or a substitution error. Reconstruction codes for single-burst-insertion/deletion were constructed in [12], where a burst of $t$ deletions/insertions means $t$ deletions or $t$ insertions occurring at consecutive positions. In these constructions, each reads allows only one type of error.
@@
In practical applications, a read may suffers from different error types, for example, both a deletion and an insertion, or both a deletion and a substitution. It was shown in [13] that a code $\mathscr{C}$ can correct $t$ deletions if and only if it can correct $t$ insertions. However, the intersection size of $t$-deletion balls of two sequences is not necessarily equal to the intersection size of their $t$-insertion balls when the intersections are not empty. Therefore, unlike the classic error correction problem, in sequence reconstruction problem, the deletion channel and the insertion channel must be treated separately. The reconstruction problem for single-insertion single-substitution was studied in [14], where the maximum intersection size of single-insertion single-substitution balls was proved to be $\lfloor\frac{n-2}{2}\rfloor\lceil\frac{n-2}{2}\rceil+4n$. The size of single-deletion multiple-substitution ball was also computed in [14], but their intersection size was not considered. In a more recent work [15], the size of the error ball for channels with multiple types of errors and at most three edits was studied. To the best of our knowledge, deriving the maximum intersection size of single-deletion single-substitution balls is still an open problem.
@@
In this work, we study the sequence reconstruction problem for $q$-ary single-deletion single-substitution channel, where $q\geq 2$ is an arbitrarily fixed integer. ((((For example, in DNA data storage, we take $q=4$.)))) We prove that for any two $q$-ary sequences with Hamming distance $d\geq 2$, the size of the intersection of their error balls is upper bounded by $2qn-3q-2-\delta_{q,2}$, where $\delta_{i,j}$ is the Kronecker delta. We also show that there exist two sequences the intersection size of whose error balls achieves this bound, which proves that the bound is tight. Note that the requirement that the Hamming distance between two sequences is at least $2$ can be satisfied by adding one parity check symbol, so results in one symbol of redundancy.
@@
The single-deletion single-substitution channel was studied in several existing works under error correction model or list-decoding model [16]--[21]. The best known single-deletion single-substitution correcting binary codes has $4\log n+o(\log n)$ bits of redundancy, where $n$ is the code length [21]. By our result, when the number of reads is $2qn-3q-1-\delta_{q,2}$, one symbol of redundancy is sufficient to guarantee correct reconstruction of the transmitted sequence.
@@
The paper is organized as follows. In Section II, we describe the problem and our main result, as well as some simple discussions that will help to prove our main result. In Section III, we develop a method for finding the intersection size of two error balls. We give a formal proof of our main result in Section IV and make conclusions and some discussions for future work in Section V.
@@ II Problem description and main result
For any integers $m\leq n$, let $[m,n]=\{m,m+1,\ldots,n\}$ and called it an interval. If $m=1$, let $[n]=[1,n]$. If $A$ is a finite set of integers, then $|A|$ is the size (cardinality) of $A$, and $\min(A)~{}($resp. $\max(A))$ is the smallest ((((resp. greatest)))) number in $A$. For simplicity, we denote $A\backslash i=A\backslash\{i\}$ for any $i\in A$. If we denote $A=\{i_{1},i_{2},\cdots,i_{k}\}$, we always assume that $i_{1}<i_{2}<\cdots<i_{k}$.
Let $\Sigma_{q}=\{0,1,\cdots,q-1\}$, where $q\geq 2$ is an arbitrarily fixed integer. For any $\bm{x}\in\Sigma_{q}^{n}$, let $x_{i}$ denote the $i$th component of $\bm{x}$ and denote $\bm{x}=x_{1}x_{2}\cdots x_{n}$ or $\bm{x}=(x_{1},x_{2},\cdots,x_{n})$. If $D=\{i_{1},i_{2},\cdots,i_{m}\}\subseteq[n]$, let $x_{D}=x_{i_{1}}x_{i_{2}}\cdots x_{i_{m}}$ and call it a subsequence of $\bm{x}$. If $D$ is an interval, $x_{D}$ is called a substring of $\bm{x}$. A run of $\bm{x}$ is a maximal substring of $\bm{x}$ consisting of identical symbols. For any two given distinct symbols $a,b\in\Sigma_{q}$, let $A_{n}(ab)$ denote the alternating sequence of length $n$ that starts with $a$ and consists of $a,b$. For example, $A_{5}(01)=01010$ and $A_{6}(01)=010101$.
For any $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$, the Hamming distance between $\bm{x}$ and $\bm{x}^{\prime}$, denoted by $d_{\text{H}}(\bm{x},\bm{x}^{\prime})$, is defined as the number of $i\in[n]$ such that $x_{i}\neq x^{\prime}_{i}$. The Levenshtein distance between $\bm{x}$ and $\bm{x}^{\prime}$, denoted by $d_{\text{L}}(\bm{x},\bm{x}^{\prime})$, is defined as the smallest integer $\ell$ such that $\bm{x},\bm{x}^{\prime}$ share some subsequence of length $n-\ell$.
Let $t$ and $s$ be non-negative integers such that $t+s<n$. For any $\bm{x}\in\Sigma_{q}^{n}$, the $t$-deletion $s$-substitution ball of $\bm{x}$, denoted by $B^{\text{D,S}}_{t,s}(\bm{x})$, is the set of all sequences that can be obtained from $\bm{x}$ by exact $t$ deletions and at most $s$ substitutions. The $t$-deletion ball of $\bm{x}$ is $B^{\text{D}}_{t}(\bm{x})\triangleq B^{\text{D,S}}_{t,0}(\bm{x})$, and the $s$-substitution ball of $\bm{x}$ is $B^{\text{S}}_{s}(\bm{x})\triangleq B^{\text{D,S}}_{0,s}(\bm{x})$. For $B\in\{B^{\text{D}}_{t},B^{\text{S}}_{s},B^{\text{D,S}}_{t,s}\}$, let
$B(\bm{x},\bm{x}^{\prime})\triangleq B(\bm{x})\cap B(\bm{x}^{\prime}).$
Given a code $\mathscr{C}\subseteq\Sigma_{q}^{n}$, let
@@
called the read coverage of $\mathscr{C}$. A central problem in sequence reconstruction is to compute $\nu(\mathscr{C};B)$ for the given $\mathscr{C}$ and $B$. Another problem is, for the given error ball function $B$ and positive integer $N$, to design a code $\mathscr{C}\subseteq\Sigma_{q}^{n}$ with $\nu(\mathscr{C};B)<N$, called an $(n,N,B)$-reconstruction code.
In this work, we assume $q\geq 2$ is a fixed positive integer. We consider the sequence reconstruction problem for $q$-ary single-deletion single-substitution channel ((((i.e., $B=B^{\text{D,S}}_{1,1})$. Our main result is the following theorem.
@@
The proof of Theorem 1 will be given in Section IV. In the rest of this section, we state some simple results that will be used by the proof.
@@ II-B Some useful observations and lemma
$B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})$ $=\bigcup_{\bm{z}\in B^{\text{D}}_{1}(\bm{x}),\bm{z}^{\prime}\in B^{\text{D}}_{1}(\bm{x}^{\prime})}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})$
$=\bigcup_{j,j^{\prime}\in[n]}B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}).$
@@
$\Lambda=\Lambda(\bm{x},\bm{x}^{\prime})$$\triangleq\Big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j,j^{\prime}\in[n]~{}\text{and}~{}$
$~{}~{}~{}~{}~{}~{}d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})\leq 2\Big{\}}.$ (2)
@@
where $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})$ and $i_{1}<i_{2}<\cdots<i_{d}$ according to our previous convention. Moreover, we let
$T^{L}=T^{L}(\bm{x},\bm{x}^{\prime})\triangleq\{i\in[2,n]:x_{i}\neq x^{\prime}_{i-1}\}.$ (4)
@@
From these definitions, it is easy to see that
$T^{R}(\bm{x},\bm{x}^{\prime})=T^{L}(\bm{x}^{\prime},\bm{x})$
We remind the reader that in the notations $T^{L}(\bm{x},\bm{x}^{\prime})$ and $T^{R}(\bm{x},\bm{x}^{\prime})$, $(\bm{x},\bm{x}^{\prime})$ is viewed as an ordered pair. Note that by the definitions, $T^{L}=T^{L}(\bm{x},\bm{x}^{\prime})\neq T^{L}(\bm{x}^{\prime},\bm{x})$ and $T^{R}=T^{R}(\bm{x},\bm{x}^{\prime})\neq T^{R}(\bm{x}^{\prime},\bm{x})$.
@@
$d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$
$=\left|\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{L}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}\right|$
@@
$d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})$
$=\left|\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{R}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}\right|$
@@
By Lemma 1, if two pairs $(j_{1},j_{2})$ and $(j^{\prime}_{1},j^{\prime}_{2})$ such that $j_{1},j_{2},j_{1}^{\prime},j_{2}^{\prime}$ satisfy the conditions of Lemma 1, then for any $(j,j^{\prime})\in[j_{1},j_{2}]\times[j_{1}^{\prime},j_{2}^{\prime}]$, we have $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=(x_{[n]\backslash j_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})=(x_{[n]\backslash j_{2}},x^{\prime}_{[n]\backslash j_{2}^{\prime}})$.

### v2

@@ Abstract
The central problem in sequence reconstruction is to find the minimum number of distinct channel outputs required to uniquely reconstruct the transmitted sequence. According to Levenshtein’s work in 2001, this number is determined by the size of the maximum intersection between the error balls of any two distinct input sequences of the channel. In this work, we study the sequence reconstruction problem for single-deletion single-substitution channel, assuming that the transmitted sequence belongs to a $q$-ary code with minimum Hamming distance at least $2$, where $q\geq 2$ is any fixed integer. Specifically, we prove that for any two $q$-ary sequences of length $n$ and with Hamming distance $d\geq 2$, the size of the intersection of their error balls is upper bounded by $2qn-3q-2-\delta_{q,2}$, where $\delta_{i,j}$ is the Kronecker delta. We also prove the tightness of this bound by constructing two sequences the intersection size of whose error balls achieves this bound.
@@ I Introduction
We consider a communication scenario where a codeword $\bm{x}$ from some codebook $\mathscr{C}$ is transmitted over a number of identical channels and the goal is to reconstruct $\bm{x}$ from all (erroneous) channel outputs (also referred to as reads in data storage applications). This problem, also known as the sequence reconstruction problem, was first proposed by Levenshtein [1], [2], and in recent years, gained more and more attentions due to its applications in DNA data storage [3]. The central problem in sequence reconstruction is to determine the minimum number of distinct channel outputs (reads) required to uniquely reconstruct $\bm{x}$. This number was shown to be equal to one plus the size of the maximum intersection between the error balls of any two distinct codewords of $\mathscr{C}$ (also referred to as the read coverage of $\mathscr{C}$ for the corresponding channel) [1]. Therefore, deriving the read coverage of $\mathscr{C}$ is critical to solving the sequence reconstruction problem. On the other hand, designing codes with given read coverage, called reconstruction codes, is also an interesting problem for sequence reconstruction.
@@
In his seminal work [1], Levenshtein studied the sequence reconstruction problem for deletion, insertion, substitution and transposition separately, where $\mathscr{C}$ is taken to be the set of all $q$-ary sequence. For the more general case that $\mathscr{C}$ is an $(\ell-1)$-deletion correcting code for some positive integer $\ell\leq t$, the problem was studied in [4] and [5] for $t$-deletion channel, and in [6] for $t$-insertion channel. Reconstruction codes for two-deletion channels can be found in [7, 8] and reconstruction codes for two-insertion channels can be found in [9]. Reconstruction codes for $q$-ary single-edit channel $(q\geq 2)$ was constructed in [10] by generalizing the construction in [11], where an edit error means a deletion, an insertion or a substitution error. Reconstruction codes for single-burst-insertion/deletion were constructed in [12], where a burst of $t$ deletions/insertions means $t$ deletions or $t$ insertions occurring at consecutive positions. In these constructions, each read is corrupted by only one type of error.
@@
In practical applications, a read may suffers from different error types, for example, both a deletion and an insertion, or both a deletion and a substitution. It was shown in [13] that a code $\mathscr{C}$ can correct $t$ deletions if and only if it can correct $t$ insertions. However, the intersection size of $t$-deletion balls of two sequences is not necessarily equal to the intersection size of their $t$-insertion balls when the intersections are not empty. Therefore, unlike the classic error correction problem, in sequence reconstruction problem, the deletion channel and the insertion channel must be treated separately. The reconstruction problem for single-insertion single-substitution was studied in [14], where the maximum intersection size of binary single-insertion single-substitution balls was proved to be $\lfloor\frac{n-2}{2}\rfloor\lceil\frac{n-2}{2}\rceil+4n$. The size of single-deletion multiple-substitution ball was also computed in [14], but their intersection size was not considered. In a more recent work [15], the size of the error ball for $q$-ary channels with multiple types of errors and at most three edits was studied. To the best of our knowledge, deriving the maximum intersection size of single-deletion single-substitution balls is still an open problem.
@@
In this work, we study the sequence reconstruction problem for $q$-ary single-deletion single-substitution channel, where $q\geq 2$ is an arbitrarily fixed integer. For example, in DNA data storage, $q$ is usually taken to be $4$. We prove that for any two $q$-ary sequences with Hamming distance $d\geq 2$, the size of the intersection of their error balls is upper bounded by $2qn-3q-2-\delta_{q,2}$, where $\delta_{i,j}$ is the Kronecker delta. We also show that there exist two sequences the intersection size of whose error balls achieves this bound, which proves that the bound is tight. Note that the requirement that the Hamming distance between two sequences is at least $2$ can be satisfied by adding one parity check symbol, so results in one symbol of redundancy.
@@
The single-deletion single-substitution channel has been studied in several existing works under the classic error correction model or list-decoding model [16]--[21]. The best known single-deletion single-substitution correcting binary codes has $4\log n+o(\log n)$ bits of redundancy, where $n$ is the code length [21]. By our result, when the number of reads is $2qn-3q-1-\delta_{q,2}$, one symbol of redundancy is sufficient to guarantee correct reconstruction of the transmitted sequence.
@@
The paper is organized as follows. In Section II, we describe the problem and our main result, as well as some simple observations that will help to prove our main result. In Section III, we develop a method for dividing the intersection of two error balls into some subsets whose size can be easily obtained. We give a formal proof of our main result in Section IV and make conclusions and some discussions for future work in Section V.
@@ II Problem description and main result
For any integers $m\leq n$, let $[m,n]=\{m,m+1,\ldots,n\}~{}($called an interval)))) and let $[n]=[1,n]$. For any set $A$, $|A|$ is the size of $A$; if $A$ is a set of numbers, then $\min(A)~{}($resp. $\max(A))$ is the smallest ((((resp. greatest)))) number in $A$. For simplicity, we denote $A\backslash i=A\backslash\{i\}$ for any $i\in A$. If we denote $A=\{i_{1},i_{2},\cdots,i_{k}\}$, we always assume that $i_{1}<i_{2}<\cdots<i_{k}$.
Let $\Sigma_{q}=\{0,1,\cdots,q-1\}$, where $q\geq 2$ is an arbitrarily fixed integer. For any $\bm{x}\in\Sigma_{q}^{n}$, let $x_{i}$ denote the $i$th component of $\bm{x}$ and write $\bm{x}=x_{1}x_{2}\cdots x_{n}$ or $\bm{x}=(x_{1},x_{2},\cdots,x_{n})$. If $D=\{i_{1},i_{2},\cdots,i_{m}\}\subseteq[n]$, let $x_{D}=x_{i_{1}}x_{i_{2}}\cdots x_{i_{m}}$ and call it a subsequence of $\bm{x}$. If $D$ is an interval, $x_{D}$ is called a substring of $\bm{x}$. A run of $\bm{x}$ is a maximal substring of $\bm{x}$ consisting of identical symbols. For any two given distinct symbols $a,b\in\Sigma_{q}$, let $A_{n}(ab)$ denote the alternating sequence of length $n$ that starts with $a$ and consists of $a,b$. For example, $A_{5}(ab)=ababa$ and $A_{6}(ab)=ababab$.
For any $\bm{x},\bm{x}^{\prime}\in\Sigma_{q}^{n}$, the Hamming distance between $\bm{x}$ and $\bm{x}^{\prime}$, denoted by $d_{\text{H}}(\bm{x},\bm{x}^{\prime})$, is defined as the number of $i\in[n]$ such that $x_{i}\neq x^{\prime}_{i}$. The Levenshtein distance between $\bm{x}$ and $\bm{x}^{\prime}$, denoted by $d_{\text{L}}(\bm{x},\bm{x}^{\prime})$, is defined as the smallest integer $\ell$ such that $\bm{x}$ and $\bm{x}^{\prime}$ share some subsequence of length $n-\ell$.
Let $t$ and $s$ be non-negative integers such that $t+s<n$. For any $\bm{x}\in\Sigma_{q}^{n}$, the $t$-deletion $s$-substitution ball of $\bm{x}$, denoted by $B^{\text{D,S}}_{t,s}(\bm{x})$, is the set of all sequences that can be obtained from $\bm{x}$ by exact $t$ deletions and at most $s$ substitutions. The $t$-deletion ball of $\bm{x}$ is $B^{\text{D}}_{t}(\bm{x})\triangleq B^{\text{D,S}}_{t,0}(\bm{x})$, and the $s$-substitution ball of $\bm{x}$ is $B^{\text{S}}_{s}(\bm{x})\triangleq B^{\text{D,S}}_{0,s}(\bm{x})$. For $B\in\{B^{\text{D}}_{t},B^{\text{S}}_{s},B^{\text{D,S}}_{t,s}\}$, let $B(\bm{x},\bm{x}^{\prime})\triangleq B(\bm{x})\cap B(\bm{x}^{\prime}).$ Given a code $\mathscr{C}\subseteq\Sigma_{q}^{n}$, let
@@
called the read coverage of $\mathscr{C}$ with respect to $B$. A central problem in sequence reconstruction is to compute $\nu(\mathscr{C};B)$, given $\mathscr{C}$ and $B$. Another problem is, given the error ball function $B$ and a positive integer $N$, to design a code $\mathscr{C}\subseteq\Sigma_{q}^{n}$ with $\nu(\mathscr{C};B)<N$, called an $(n,N,B)$-reconstruction code.
In this work, we assume $q\geq 2$ is any fixed positive integer and consider the sequence reconstruction problem for $q$-ary single-deletion single-substitution channel ((((i.e., $B=B^{\text{D,S}}_{1,1})$. Our main result is the following theorem.
@@
The proof of Theorem 1 will be given in Section IV. In the rest of this section, we state some simple observations that will be used in our proof.
@@ II-B Some useful observations and lemma
$B^{\text{D,S}}_{1,1}(\bm{x},\bm{x}^{\prime})=\bigcup_{\bm{z}\in B^{\text{D}}_{1}(\bm{x}),\bm{z}^{\prime}\in B^{\text{D}}_{1}(\bm{x}^{\prime})}B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})=\bigcup_{j,j^{\prime}\in[n]}B^{\text{S}}_{1}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}).$
@@
$\Lambda=\Lambda(\bm{x},\bm{x}^{\prime})\triangleq\Big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j,j^{\prime}\in[n]~{}\text{and}~{}d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})\leq 2\Big{\}}.$ (2)
@@
where $d=d_{\text{H}}(\bm{x},\bm{x}^{\prime})$ and $i_{1}<i_{2}<\cdots<i_{d}$ according to our previous convention. We further let
$T^{L}=T^{L}(\bm{x},\bm{x}^{\prime})\triangleq\{i\in[2,n]:x_{i}\neq x^{\prime}_{i-1}\}$ (4)
@@
From these definitions, it is easy to see that $T^{R}(\bm{x},\bm{x}^{\prime})=T^{L}(\bm{x}^{\prime},\bm{x})$. Note that in the notations $T^{L}(\bm{x},\bm{x}^{\prime})$ and $T^{R}(\bm{x},\bm{x}^{\prime})$, $(\bm{x},\bm{x}^{\prime})$ is viewed as an ordered pair. Moreover, by the definitions, $T^{L}=T^{L}(\bm{x},\bm{x}^{\prime})\neq T^{L}(\bm{x}^{\prime},\bm{x})$ and $T^{R}=T^{R}(\bm{x},\bm{x}^{\prime})\neq T^{R}(\bm{x}^{\prime},\bm{x})$.
@@
$d_{\text{H}}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})$ $=\left|\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{L}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}\right|$
@@
$d_{\text{H}}(x_{[n]\backslash j^{\prime}},x^{\prime}_{[n]\backslash j})$ $=\left|\big{(}S\cap[1,j-1]\big{)}\cup\big{(}T^{R}\cap[j+1,j^{\prime}]\big{)}\cup\big{(}S\cap[j^{\prime}+1,n]\big{)}\right|$
@@
By Lemma 1, if $j_{1},j_{2},j_{1}^{\prime},j_{2}^{\prime}$ satisfy the conditions of Lemma 1, then for any $(j,j^{\prime})\in[j_{1},j_{2}]\times[j_{1}^{\prime},j_{2}^{\prime}]$, we have $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})=(x_{[n]\backslash j_{\ell}},x^{\prime}_{[n]\backslash j_{\ell^{\prime}}^{\prime}})$ for any ,$\ell,\ell^{\prime}\in\{1,2\}$.